Consider a system running 200 I/O-bound jobs and 10 CPU-bound job. Assume that the I/O-bound jobs issue an I/O operation once for every millisecond of CPU computing and that each I/O operation takes 10 milliseconds to complete. Also assume that the context switching overhead is 0.1millisecond and that all processes are long-running tasks. What percentage of CPU time is wasted in context switch for a round-robin scheduler when:

(a) The time quantum is 1 milliseconds

(b) The time quantum is 10 milliseconds
Answer #1

Answer:

(a) The time quantum is 1 millisecond:

Irrespective of which process is scheduled, the scheduler incurs a 0.1 millisecond context-switching cost for every context-switch.

Therefore, percentage of CPU time is wasted in context switch = (1/1.1) * 100 = 90.90%.

(b) The time quantum is 10 milliseconds:

The I/O-bound tasks incur a context switch after using up only 1 millisecond of the time quantum.

The time required to cycle through all the processes is therefore (1ms + 0.1ms) ∗ 200 + (10ms + 0.1ms) * 10 = 321ms

(as each I/O-bound task executes for 1 millisecond and then incur the context switch task, whereas the CPU-bound task executes for 10 milliseconds before incurring a context switch).

Actual total time taken via both the jobs = 1 * 200 + 10 * 10 = 300

Therefore, percentage of CPU time is wasted in context switch = (300 / 321 ) * 100 = 93.45%

The CPU utilization is therefore 20/21.1 * 100 = 94%.

Please give thumbsup, if you like it. Thanks.